# GENIE Collaboration

### v1

@@ III.1 Data Observables
Table 3: Observables of the TKI measurements and their binning. Those with $\checkmark$”’s are used for tuning, while those without are for the respective validation. See later text for the definitions of Combi-Best-RedPar and Combi-Best-AllPar.
Figure 4: Change of $\chi^{2}$ calculated for the full (i.e., tuned plus validation) observable set as a function of the tuned combination (cf. Table 6 in Appendix .1). The two model parameter sets (AllPar and RedPar, see later text for definitions) are compared and it can be seen that the respective minima happen at Combi-15 and 26.
III.2 Model Parameters
As outlined in Sec. II, we focus exclusively on the currently tunable parameters within the SF-LFG and hA models. Therefore, $14$ parameters (collectively referred to as the AllPar set) are utilized for tuning, as detailed in Table 4. Nominal values and associated uncertainties for the hA model are taken from Table 17.3 in Ref. [49].
@@ III.2 Model Parameters
Table 4: Tuneable parameters and their ranges in the SF-LFG (uppermost group) and hA (lower groups, uncertainties from Ref. [49]) models. As the SF-LFG model has not been tuned, no official uncertainties are available. Parameters to be tuned in the two sets are marked with “✓”’s. See later text for definitions of G24-t and G24-c.
In our analysis of the SF-LFG model, the two most relevant parameters are $R_{\textrm{SRC}}$, representing the fraction of the high-momentum tail from short-range correlations (SRC), notably above the Fermi surface, and $E_{\textrm{RM}}^{\textrm{C}}$, marking the commencement of the nuclear removal energy distribution for carbon. A larger $R_{\textrm{SRC}}$ indicates the presence of more energetic initial nucleons, while an increased $E_{\textrm{RM}}^{\textrm{C}}$ implies that greater energy is necessary to liberate a nucleon from the carbon nucleus, so the product particles will possess lower energy. Given the novelty and limited constraints of the spectral-function-like implementation, this study employs relatively relaxed priors for $R_{\textrm{SRC}}$ and $E_{\textrm{RM}}^{\textrm{C}}$, at 0.12$\pm$0.12 and 0.01$\pm$0.005 GeV, respectively.
genie employs the hA model [49] for the majority of neutrino analysis, owing to its straightforward reweighting capability. Unlike cascade models such as the hN model [49], this model determines the FSI type for each hadron exactly once, without considering further propagation of rescattered hadrons within the nucleus. Its model parameters are explained as follows.
@@
    where $\sigma_{\textrm{hN,tot}}$ is the total hadron-nucleon cross section and $\rho$ is the number density of the nucleons inside the nucleus. Once a hadron is generated inside the nucleus, it travels a distance $\lambda$ before a probabilistic determination of rescattering occurs, inversely related to $\lambda$. If no rescattering happens, the hadron will be propagated for another $\lambda$ and the dice will be thrown again. The cycle repeats until a rescattering happens or the hadron is propagated far enough to escape the nucleus without any interactions. If a rescattering happens, a dice is thrown again to determine the type of rescattering. The relative probabilities of each type are extracted from stored hadron scattering data [54, 55, 56, 57, 58, 59, 60]. After a specific rescattering type is chosen, the rescattered product will be generated. These rescattered products will not undergo further rescattering and will be transported out of the nucleus as final particles. The MFP values for pions and nucleons can be modified using scaling factors ($S_{\lambda}^{\pi^{\pm}}$, $S_{\lambda}^{\pi^{0}}$, and $S_{\lambda}^{\textrm{N}}$) detailed in Table 4. Both changed pions, $\pi^{+}$ and $\pi^{-}$, are scaled by the same factor, $S_{\lambda}^{\pi^{\pm}}$, due to isospin symmetry. The $\lambda$ for $\pi^{0}$ is calculated from those of charged pions based on isospin symmetry as well and can be adjusted separately by $S_{\lambda}^{\pi^{0}}$. If the factor is larger than $1$, $\lambda$ will increase and hence the total number of steps required for the hadron to escape the nucleus reduces, thereby reducing the average probability of rescattering. Given the high precision with which total hadron-nucleus cross sections are determined [54, 55, 56, 57, 58], strict Gaussian priors (error—$\sigma$, not to be confused with the various cross sections—of $0.2$, same as the systematic uncertainties shown in Table 4) are applied to the MFP scaling factors. Varying $S_{\lambda}^{\pi^{0}}$ by ±$\pm 2.5~{}\sigma$ significantly affects the MINERvA-$\pi^{0}$ $\textrm{d}\sigma/\textrm{d}p_{\textrm{N}}$, as depicted in Fig. 5. Decreasing MFP naturally leads to increased rescattering. Hence, fewer $\pi^{0}$s can escape the nucleus and fewer events will have $\pi^{0}$ in the final states, thereby reducing the cross section.
@@
    The relative probability of each rescattering type can be adjusted by a scale factor. There are four rescattering types for both nucleons and pions available for tuning. They are charge exchange (CEX), inelastic scattering (INEL), absorption (ABS), and pion production (PIPD). The default energy-dependent probability distributions come from hadron data tuning [54, 55, 56, 57, 58] and can be adjusted by respective scaling factors, like $S_{\textrm{CEX}}^{\pi}$ for CEX of all pions, listed in Table 4. The relative probabilities of all four rescattering type will be normalized such that they sum up to one. Scaling factors directly multiply the probability of each type, with new relative probabilities subsequently renormalized. More specifically, suppose the initial fraction of absorption, $f_{\textrm{ABS}}$, is $0.5$, which means that the sum of the other fractions is $f_{\textrm{other}}=f_{\textrm{INEL}}+f_{\textrm{CEX}}+f_{\textrm{PIPD}}=1-0.5=0.5$. If $f_{\textrm{ABS}}$ is scaled up with $S_{\textrm{ABS}}=2$, the new absorption fraction is
@@
    which is not equal to simply multiplying $f_{\textrm{ABS}}$ by $S_{\textrm{ABS}}$ due to the presence of the other rescattering types. Hence, it is unavoidable that a change of the scaling factor for one type would affect others. Nonetheless, when the cross section is plotted with a breakdown according to rescattering type, interpretation will be clear regarding the increase or decrease of a particular FSI type. This correlation implies that the scaling parameters’ effective impact on FSI fate is often less pronounced than what the raw numbers suggest. As illustrated in the above example, $f_{\textrm{ABS}}$ is only scaled up by $f^{\prime}_{\textrm{ABS}}/f_{\textrm{ABS}}=4/3=1.3$ instead of $S_{\textrm{ABS}}=2$. Hence, a relatively more relaxed Gaussian prior, with a $\sigma$ of $0.5$, is placed on the FSI fate scales, slightly larger than the systematic uncertainties shown in Table 4.
@@
IV The First TKI-Driven GENIE Tunes
Figure 6: Correlation coefficient for AllPar on Combi-26.
Figure 7: Similar to Fig. 6 but for RedPar.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 8: Similar to Fig. 2 but with G24-c. The nominal tune, G24-0, is also plotted for comparison.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 9: Similar to Fig. 9 but for the $p_{\textrm{N}}$ ($\delta p_{\textrm{T}}$ for T2K-$0\pi$) measurements.
Given the intricate correlations among model parameters, determining which ones are most sensitive to the data proves challenging. Considerable correlation and anti-correlation are observed between different FSI fates for the same particle, e.g. $S_{\textrm{ABS}}^{\pi^{\pm}}$ and $S_{\textrm{CEX}}^{\pi}$, as shown in Figs. 6 and 7. Our tuning began with the AllPar parameter set, achieving the best result with Combi-15, which is denoted as Combi-Best-AllPar in Table 3 (cf. Fig. 4 for the $\chi^{2}$ distribution of all combinations). Observations showed that for most combinations, certain parameters remained close to their default values, within $1$ $\sigma$ of the imposed priors, such as $S_{\textrm{PIPD}}^{\pi}$ since none of the data sets has a significant contribution from PIPD of pions—Removing these parameters from the tuning process does not significantly affect the outcome. Therefore, we then constructed a reduced set comprising only $R_{\textrm{SRC}}$, $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{CEX}}^{\textrm{N}}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$ (denoted as RedPar in Table 4) and ran the tuning on the 26 combinations with it. Note that the parameters most correlated with $S_{\textrm{CEX}}^{\pi}$ and $S_{\lambda}^{\pi^{0}}$ are not present in RedPar. The RedPar tuning proved more stable, with nearly all combinations showing a more negative $\chi^{2}$ change than the AllPar tuning as shown in Fig. 4. Moreover, the single-observable (Combi-1 to 5), single-measurement (Combi-6 to 9), single-experiment (Combi-10 and 11), and single-topology (Combi-12 and 13) combinations all systematically yielded no or very limited overall (tuned plus validation) improvement post-tuning. The best tuning then occurred with Combi-26 (Combi-Best-RedPar in Table 3), that is, all TKI variables excluding $\delta\phi_{\textrm{T}}$ and $\delta p_{\textrm{T}}$ (unless no $p_{\textrm{N}}$ is available)—$\delta\alpha_{\textrm{T}}$, $p_{\textrm{N}}$ (or $\delta p_{\textrm{T}}$ if $p_{\textrm{N}}$ is unavailable), and $\delta p_{\textrm{TT}}$—from pionless and pion-production measurements in both T2K and MINERvA.
Table 5 summarizes the primary results of our tuning process. The tuned RedPar, G24_20i_06_22c (G24-c), is derived from the observable set Combi-Best-RedPar, and the tuned AllPar, G24_20i_fullpara_alt (G24-t), is from Combi-Best-AllPar. The upper part of Table 5 contains the parameter values. For the G24-c tune, changes to the SF-LFG model were moderately sized ($R_{\textrm{SRC}}$ increases from $0.12$ to $0.15$), while in the hA model, $S_{\lambda}^{\pi^{0}}$ and $S_{\textrm{CEX}}^{\pi}$ are highly suppressed and the nucleon FSI has larger CEX and PIPD and lower ABS. The tune G24-t differs most significantly from G24-c in $R_{\textrm{SRC}}$ and $S_{\textrm{CEX}}^{\pi}$. The table’s lower section illustrates $\chi^{2}$ improvements for selected observables and their corresponding validation sets. Crucially, tuning should also enhance the accuracy of validation set descriptions. Overall, both tunes have substantial reduction in the total $\chi^{2}$. In the following, we will discuss the two tunes.
@@ IV The First TKI-Driven GENIE Tunes
Table 5: Parameters in G24-0, G24-c, and G24-t. Those without errors are not tuned. Lower section: combi indicates that the following $\chi^{2}$ sums are calculated for the tuned measurements (i.e. Combi-Best-RedPar and Combi-Best-AllPar for G24-c and G24-t, respectively), while vald indicates that the respective validation sets are used; untuned means that the $\chi^{2}$ calculation uses the nominal values of the parameters, while tuned denotes the tuned ones. diff displays the improvement achieved by the respective tuning.
IV.1 Reduced Tune: G24-c
@@
In MINERvA-$\pi^{0}$, the number of $\pi^{0}$s undergoing no FSI (“No Interaction” as shown in the figures) is adjustable via the $S_{\lambda}^{\pi^{0}}$ parameter. A decrease in $S_{\lambda}^{\pi^{0}}$ reduces the size of the “No Interaction” events, as is shown in Fig. 10b when compared to Fig. 10a. Meanwhile, an increase in $\pi^{0}$ rescattering only manifests in pion-less measurements through ABS. There is indeed an increase of ABS for $\pi^{0}$ in T2K-$0\pi$ and MINERvA-$0\pi$, but the fraction is small to begin with, so the impact is small. The increase of $\pi^{0}$ rescattering can increase T2K-$\pi^{+}$ via CEX as discussed below. However, due to the significant suppression of CEX, its impact on T2K-$\pi^{+}$ is also minimal (for detailed breakdown, see Figs. 14-16 in Appendix .2).
As shown in Fig. 10a, the MINERvA-$\pi^{0}$ prediction from the nominal tune has considerable contributions from CEX controlled by $S_{\textrm{CEX}}^{\pi}$. It does not impact T2K-$0\pi$ and MINERvA-$0\pi$ as CEX only changes the pion type without removing them. Hence, events with pions in the final state would be rejected in these two data sets regardless of the pion charge. In principle, CEX will also migrate events between the signal and background definitions for T2K-$\pi^{+}$ when a $\pi^{0}$ is converted to a $\pi^{+}$ and vice verse. However, T2K-$\pi^{+}$ does not have a considerable CEX fraction to begin with, so changing CEX will have a minimal impact on its prediction. Hence, this is one of the most effective paths to decrease the MINERvA-$\pi^{0}$ cross section prediction without affecting other measurements, and indeed the $S_{\textrm{CEX}}^{\pi}$ is heavily suppressed as shown in Table 5 and illustrated in Fig. 10b compared to Fig. 10a. However, due to the intricate correlation between the FSI fates in the hA implementation discussed in Sec. III.2, although $S_{\textrm{INEL}}^{\pi}$ and $S_{\textrm{PIPD}}^{\pi}$ are not modified explicitly, a considerable increase is observed in both INEL and PIPD for MINERvA-$\pi^{0}$.
@@
Refer to captionRefer to caption
Figure 11: Similar to Fig. 10 but for $p_{\textrm{N}}$ with proton FSI fate decomposition. Note that the “Absorption” of the proton is referring to the absorption of the $\pi^{+}$ from the decay of $\Delta^{++}$, which could lead to emission of nucleons as discussed in Sec. III.2.
IV.2 Alternative Tune: G24-t

### v2

@@ 4.1 Data points
Table 3: Observables of the TKI measurements and their binning. Those with $\checkmark$”’s are used for tuning, while those without are for the respective validation. See Table 4 for definitions of Combi-Best-RedPar and Combi-Best-AllPar.
Figure 5: Change of $\chi^{2}$ calculated for the full (i.e., tuned plus validation) observable set as a function of the tuned combination (cf. Table 6 in Appendix A). The two model parameter sets (AllPar and RedPar, see Table 4 for definitions) are compared and it can be seen that the respective minima happen at Combi-15 and 26.
4.2 Model parameters
As outlined in Sec. 3, we focus exclusively on the currently tunable parameters within the SF-CFG and hA models. Therefore, $14$ parameters (collectively referred to as the AllPar set) are utilized for tuning, as detailed in Table 4. Nominal values and associated uncertainties for the hA model are taken from Table 17.3 in Ref. Andreopoulos:2015wxa .
@@ 4.2 Model parameters
Table 4: Tuneable parameters and their ranges in the SF-CFG (uppermost group) and hA (lower groups, uncertainties from Ref. Andreopoulos:2015wxa ) models. Parameters to be tuned in the two sets are marked with “✓”’s. See later text for definitions of G24-t and G24-c.
In our analysis of the SF-CFG model, the two most relevant parameters are $R_{\textrm{SRC}}$, representing the fraction of the high-momentum tail from short-range correlations (SRC), notably above the Fermi surface, and $E_{\textrm{RM}}^{\textrm{C}}$, marking the commencement of the nuclear removal energy distribution for carbon. A larger $R_{\textrm{SRC}}$ indicates the presence of more energetic initial nucleons, while an increased $E_{\textrm{RM}}^{\textrm{C}}$ implies that greater energy is necessary to liberate a nucleon from the carbon nucleus, so the product particles will possess lower energy. Given the novelty and limited constraints of the spectral-function-like implementation, this study employs relatively relaxed priors for $R_{\textrm{SRC}}$ and $E_{\textrm{RM}}^{\textrm{C}}$, at 0.12$\pm$0.12 and 0.01$\pm$0.005 GeV, respectively.
genie users employ the hA model Andreopoulos:2015wxa for the majority of neutrino analysis, owing to its straightforward reweighting capability. Unlike cascade models such as the hN model Andreopoulos:2015wxa , this model determines the FSI type for each hadron exactly once, without considering further propagation of rescattered hadrons within the nucleus. Its model parameters are explained as follows.
@@
    where $\sigma_{\textrm{hN,tot}}$ is the total hadron-nucleon cross section and $\rho$ is the number density of the nucleons inside the nucleus. Once a hadron is generated inside the nucleus, it travels a distance $\lambda$ before a probabilistic determination of rescattering occurs, inversely related to $\lambda$. If no rescattering happens, the hadron will be propagated for another $\lambda$ and the dice will be thrown again. The cycle repeats until a rescattering happens or the hadron is propagated far enough to escape the nucleus without any interactions. If a rescattering happens, the type of rescattering is evaluated according to their cross section. The relative probabilities of each type are extracted from stored hadron scattering data LADS:1999dyv ; Navon:1983xj ; Carroll:1976hj ; Clough:1974qt ; BAUHOFF1986429 ; Mashnik:2000up ; Ishibashi:1997gbe . After a specific rescattering type is chosen, the rescattered product will be generated. These rescattered products will not undergo further rescattering and will be considered out of the nucleus as final state particles. The MFP values for pions and nucleons can be modified using scaling factors ($S_{\lambda}^{\pi^{\pm}}$, $S_{\lambda}^{\pi^{0}}$, and $S_{\lambda}^{\textrm{N}}$) detailed in Table 4. Both changed pions, $\pi^{+}$ and $\pi^{-}$, are scaled by the same factor, $S_{\lambda}^{\pi^{\pm}}$, due to isospin symmetry. The $\lambda$ for $\pi^{0}$ is calculated from those of charged pions based on isospin symmetry as well and can be adjusted separately by $S_{\lambda}^{\pi^{0}}$. If the factor is larger than $1$, $\lambda$ will increase and hence the total number of steps required for the hadron to escape the nucleus reduces, thereby reducing the average probability of rescattering. Given the high precision with which total hadron-nucleus cross sections are determined LADS:1999dyv ; Navon:1983xj ; Carroll:1976hj ; Clough:1974qt ; BAUHOFF1986429 , strict Gaussian priors (error—$\sigma$, not to be confused with the various cross sections—of $0.2$, same as the systematic uncertainties shown in Table 4) are applied to the MFP scaling factors. Varying $S_{\lambda}^{\pi^{0}}$ by ±$\pm 2.5~{}\sigma$ significantly affects the MINERvA-$\pi^{0}$ $\textrm{d}\sigma/\textrm{d}p_{\textrm{N}}$, as depicted in Fig. 6. Decreasing MFP naturally leads to increased rescattering. Hence, fewer $\pi^{0}$s can escape the nucleus and fewer events will have $\pi^{0}$ in the final states, thereby reducing the cross section.
@@
    The relative probability of each rescattering type can be adjusted by a scale factor. There are four rescattering types for both nucleons and pions available for tuning. They are charge exchange (CEX), inelastic scattering (INEL), absorption (ABS), and pion production (PIPD). The default energy-dependent probability distributions come from hadron data tuning LADS:1999dyv ; Navon:1983xj ; Carroll:1976hj ; Clough:1974qt ; BAUHOFF1986429 and can be adjusted by respective scaling factors, like $S_{\textrm{CEX}}^{\pi}$ for CEX of all pions, listed in Table 4. The relative probabilities of all four rescattering types will be normalized such that they sum up to one. Scaling factors directly multiply the probability of each type, with new relative probabilities subsequently renormalized. More specifically, suppose the initial fraction of absorption, $f_{\textrm{ABS}}$, is $0.5$, which means that the sum of the other fractions is $f_{\textrm{other}}=f_{\textrm{INEL}}+f_{\textrm{CEX}}+f_{\textrm{PIPD}}=1-0.5=0.5$. If $f_{\textrm{ABS}}$ is scaled up with $S_{\textrm{ABS}}=2$, the new absorption fraction is
@@
    which is not equal to simply multiplying $f_{\textrm{ABS}}$ by $S_{\textrm{ABS}}$ due to the presence of the other rescattering types. Hence, it is unavoidable that a change of the scaling factor for one type would affect others. Nonetheless, when the cross section is plotted with a breakdown according to rescattering type, the interpretation will be clear regarding the increase or decrease of a particular FSI type. This correlation implies that the scaling parameters’ effective impact on FSI fate is often less pronounced than what the raw numbers suggest. As illustrated in the above example, $f_{\textrm{ABS}}$ is only scaled up by $f^{\prime}_{\textrm{ABS}}/f_{\textrm{ABS}}=4/3=1.3$ instead of $S_{\textrm{ABS}}=2$. Hence, a relatively more relaxed Gaussian prior, with a $\sigma$ of $0.5$, is placed on the FSI fate scales, slightly larger than the systematic uncertainties shown in Table 4.
@@
5 The first TKI-driven genie tunes
Given the intricate correlations among model parameters, determining which ones are most sensitive to the data proves challenging. Considerable correlation and anti-correlation are observed between different FSI fates for the same particle, e.g. $S_{\textrm{ABS}}^{\pi^{\pm}}$ and $S_{\textrm{CEX}}^{\pi}$, as shown in Figs. 7(a) and 7(b). Our tuning began with the AllPar parameter set, achieving the best result with Combi-15, which is denoted as Combi-Best-AllPar in Table 3 (cf. Fig. 5 for the $\chi^{2}$ distribution of all combinations). Observations showed that for most combinations, certain parameters remained close to their default values, within $1$ $\sigma$ of the imposed priors, such as $S_{\textrm{PIPD}}^{\pi}$ since none of the data sets has a significant contribution from PIPD of pions—Removing these parameters from the tuning process does not significantly affect the outcome. Therefore, we then constructed a reduced set comprising only $R_{\textrm{SRC}}$, $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{CEX}}^{\textrm{N}}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$ (denoted as RedPar in Table 4) and ran the tuning on the 26 combinations with it. Note that the parameters most correlated with $S_{\textrm{CEX}}^{\pi}$ and $S_{\lambda}^{\pi^{0}}$ are not present in RedPar. The RedPar tuning proved more stable, with nearly all combinations showing a more negative $\chi^{2}$ change than the AllPar tuning, as shown in Fig. 5. Moreover, the single-observable (Combi-1 to 5), single-measurement (Combi-6 to 9), single-experiment (Combi-10 and 11), and single-topology (Combi-12 and 13) combinations all systematically yielded no or limited overall (tuned plus validation) improvement post-tuning. The best tuning then occurred with Combi-26 (Combi-Best-RedPar in Table 3), that is, all TKI variables excluding $\delta\phi_{\textrm{T}}$ and $\delta p_{\textrm{T}}$ (unless no $p_{\textrm{N}}$ is available)—$\delta\alpha_{\textrm{T}}$, $p_{\textrm{N}}$ (or $\delta p_{\textrm{T}}$ if $p_{\textrm{N}}$ is unavailable), and $\delta p_{\textrm{TT}}$—from pionless and pion-production measurements in both T2K and MINERvA.
(a) AllPar
(b) RedPar
Figure 7: Post-fit Correlation coefficient on Combi-26.
@@ 5 The first TKI-driven genie tunes
Table 5: Parameters in G24-0, G24-c, and G24-t. Those without errors are not tuned. Lower section: combi indicates that the following $\chi^{2}$ sums are calculated for the tuned measurements (i.e. Combi-Best-RedPar (98 bins) and Combi-Best-AllPar (72 bins) for G24-c and G24-t, respectively), while vald indicates that the respective validation sets are used; untuned means that the $\chi^{2}$ calculation uses the nominal values of the parameters, while tuned denotes the tuned ones. diff displays the improvement achieved by the respective tuning.
Table 5 summarizes the primary results of our tuning process. The tuned RedPar, G24_20i_06_22c (G24-c), is derived from the observable set Combi-Best-RedPar, and the tuned AllPar, G24_20i_fullpara_alt (G24-t), is from Combi-Best-AllPar. The upper part of Table 5 contains the parameter values. For the G24-c tune, changes to the SF-CFG model were moderately sized ($R_{\textrm{SRC}}$ increases from $0.12$ to $0.15$), while in the hA model, $S_{\lambda}^{\pi^{0}}$ and $S_{\textrm{CEX}}^{\pi}$ are highly suppressed and the nucleon FSI has larger CEX and PIPD and lower ABS. The tune G24-t differs most significantly from G24-c in $R_{\textrm{SRC}}$ and $S_{\textrm{CEX}}^{\pi}$. The table’s lower section illustrates $\chi^{2}$ improvements for selected observables and their corresponding validation sets. Crucially, tuning should also enhance the accuracy of validation set descriptions. Overall, both tunes have substantial reduction in the total $\chi^{2}$. In the following, we will discuss the two tunes.
5.1 Reduced tune: G24-c
@@
Figure 8: Similar to Fig. 3 but with G24-c. The G24-0 prediction is also plotted for comparison.
Figure 9: Similar to Fig. 9 but for the $p_{\textrm{N}}$ (T2K-$\pi^{+}$, MINERvA-$0\pi$ and MINERvA-$\pi^{0}$) and $\delta p_{\textrm{T}}$ (T2K-$0\pi$) measurements.
@@
In MINERvA-$\pi^{0}$, the number of $\pi^{0}$s undergoing no FSI (“No Interaction”, as shown in the figures) is adjustable via the $S_{\lambda}^{\pi^{0}}$ parameter. A decrease in $S_{\lambda}^{\pi^{0}}$ reduces the size of the “No Interaction” events, as is shown in Fig. 10b when compared to Fig. 10a. Meanwhile, an increase in $\pi^{0}$ rescattering only manifests in pion-less measurements through ABS. There is indeed an increase of ABS for $\pi^{0}$ in T2K-$0\pi$ and MINERvA-$0\pi$, but the fraction is small to begin with, so the impact is small. The increase of $\pi^{0}$ rescattering can increase T2K-$\pi^{+}$ via CEX as discussed below. However, due to the significant suppression of CEX, its impact on T2K-$\pi^{+}$ is also minimal (for detailed breakdown, see Figs. 15-17 in Appendix B).
As shown in Fig. 10a, the MINERvA-$\pi^{0}$ prediction from the nominal tune has considerable contributions from CEX controlled by $S_{\textrm{CEX}}^{\pi}$. It does not impact T2K-$0\pi$ and MINERvA-$0\pi$ as CEX only changes the pion type without removing them. Hence, events with pions in the final state would be rejected in these two data sets regardless of the pion charge. In principle, CEX will also migrate events between the signal and background definitions for T2K-$\pi^{+}$ when a $\pi^{0}$ is converted to a $\pi^{+}$ and vice verse. However, T2K-$\pi^{+}$ does not have a considerable CEX fraction to begin with, so changing CEX will have a minimal impact on its prediction. Hence, this is one of the most effective paths to decrease the MINERvA-$\pi^{0}$ cross section prediction without affecting other measurements, and indeed the $S_{\textrm{CEX}}^{\pi}$ is heavily suppressed as shown in Table 5 and illustrated in Fig. 10b compared to Fig. 10a. However, due to the intricate correlation between the FSI fates in the hA implementation discussed in Sec. 4.2, although $S_{\textrm{INEL}}^{\pi}$ and $S_{\textrm{PIPD}}^{\pi}$ are not modified explicitly, a considerable increase is observed in both INEL and PIPD for MINERvA-$\pi^{0}$.
Figure 11: Similar to Fig. 10 but for $p_{\textrm{N}}$ with proton FSI fate decomposition. Note that the “Absorption” of the proton is referring to the absorption of the $\pi^{+}$ from the decay of $\Delta^{++}$, which could lead to emission of nucleons as discussed in Sec. 4.2.
@@
5.2 Alternative tune: G24-t